• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 21:57:32 on June16,2009

    Entry number 281811

    This entry is a followup to: 281789

    keyword=Compton activity, part III

    Edet hit plane distributions look bizarre -- big background
    that drops off from beam position -- no clear Compton edge.

    Slits were fully open, the MCC had reduced laser power.
    Will increase laser power, reduce slits to see if the
    background improves.

    21:38 beam returns. 3 uA, slits not reduced yet
    Run 17554

    Run 17555 -- mixed conditions (ignore)

    21:48 closed slits down. Run 17556. Threshold = 30.
    Hit distributions in fig 1.

    21:50 stop run. move detector out (to 60 mm)

    Will resume high-current running for PVDIS while
    we think about what to do -- it seems the background
    is enormous at present and there is no hint of
    Compton signal.

    It might help to trigger on photons with standard
    firmware and/or try Michel's firmware which triggers
    on photons -- we'll discuss it.




    FIGURE 1